Do the following:

A. If

= arcsin

then what is sin(theta)?

B. Draw the reference triangle for

= arcsin

and find all the sides.

C. Manually find the exact value of the expression

sec (2 arcsin :-)
Answer #1

son we have in A L o = sin(4) I sino sin(sin (4) sino = 1/2 B) y sino = 1 = 1 41348 |0o= (1? + B2 | 49 = 1 +B2 Ą 8= 49-12/488420 1-2 sin²o 1-2 so, sec (2 sint 1) = 49